Calculate the dollar-based percentage return an American would have if he bought a British stock at ₤50 per share and sold it one year later at ₤60. The spot exchange rate one year ago was $1.50 = ₤1 and the spot rate prevailing at the end of the year was $1.20 = ₤1.
Answer #1

Stock bought at ₤50 per share

So,

Dollar Price = 1.50(50) = $75

Stock sold at ₤60

So,

Dollar Selling Price = 1.20(60) = $72

Dollar Percentage Return = (72 - 75)/75

Dollar Percentage Return = -4%